Let D={(x,y)∈R^2:y^2<x<y and 0<y<1}. The answer to ∬D 2y^2xdA

is
Select one:
a. 70
b. 1/35
c. -70
d. 2/35
e. 35
1
Expert's answer
2020-06-02T18:25:03-0400

Since region bounded by

"x:y^2\to y,\ \ \ \ \ \ \ \ y :0\to 1"

Then

"\begin{aligned} \int\int_{D} y^2xdA&= \int_{0}^{1}\int_{y^2}^{y} y^2xdxdy \\
&= \frac{1}{2}\int_{0}^{1} y^2x^2\bigg|_{y^2}^{y} dy \\
&=\frac{1}{2}\int_{0}^{1} y^2(y^2-y^4)dy\\
&= \frac{1}{2}\int_{0}^{1} (y^4-y^6)dy\\
&= \frac{1}{2} (\frac{1}{5}y^5-\frac{1}{7}y^7)\bigg|_{0}^{1}\\
&= \frac{1}{2}( \frac{1}{5} -\frac{1}{7})\\
&= \frac{1}{35}
 \end{aligned}"